OpenStudy (anonymous):

Express the following in standard notation: 2 x 10^-5 A. 200,000 B. 20,000 C. 0.00002 D. 0.000002

OpenStudy (anonymous):

c

OpenStudy (anonymous):

\[2\times 10^-5 = \frac{2}{10^5}\]

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

I'm a little confused.. which one would it be then?

OpenStudy (anonymous):

Well if you do the math \[\frac{2}{100000} = 0.00002

OpenStudy (anonymous):

If you see carefully you can note that there is 5 ceros at the left of the number 2.

OpenStudy (anonymous):

ohhhh ! i see now ! thanks (:

OpenStudy (anonymous):

You're welcome =).

OpenStudy (anonymous):

(: